OpenStudy (anonymous):

Suppose a forest fire spreads in a circle with radius at rate of 5 per minute. when the radius reaches 200 feet, at what rate is the area of the burning region increasing?

OpenStudy (anonymous):

\[\frac{dr}{dt}=5\]\[\frac{dA}{dt}=2 \pi r\frac{dr}{dt}=10 \pi r\]For r=200, the rate is 2000pi

OpenStudy (anonymous):

woow same as my asnwer :3 .. thanks man im giving u a medal ; ]